OpenStudy (anonymous):

solve equation by graphing y=-x-8 9x-8y=-21 please someone help me step by step so i can learn to do this on my own

OpenStudy (amistre64):

do you have graph paper?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

then establish a horizontal line and call it X; draw a vertical line and call it Y

OpenStudy (anonymous):

did that

OpenStudy (amistre64):

where the 2 lines cross, call it (0,0) and we can then plot other points from that origin.

OpenStudy (amistre64):

get 2 different colored crayons , one for each line that we are going to plot...

OpenStudy (anonymous):

i have numbered graph paper

OpenStudy (anonymous):

k

OpenStudy (amistre64):

numbered what???....... thats just cheating ;) y=-x-8 since this is an equation for a line, we only need 2 points to draw a line with right?

OpenStudy (anonymous):

what, no my graph paper is pre-numbered y,x 10,-10

OpenStudy (amistre64):

back in my day we had to work to get numbers on graph paper lol

OpenStudy (anonymous):

yeah i was doing that way till i found it online and printed a bunch

OpenStudy (amistre64):

online???...... oh well :) So lets plot 2 points for this y=-x-8 equation. What are your 2 favorite numbers for an x ?

OpenStudy (anonymous):

any #'s?

OpenStudy (amistre64):

any number between -10 and 10 would be helpful :)

OpenStudy (anonymous):

5, -3

OpenStudy (anonymous):

he easy form is must include : -name of a drug -history of drug -street name -short +long term effect -how is it ingested -what effects does it have on body -anyother intersting facts amisstro plz help me write an essay

OpenStudy (amistre64):

ok...les start with ....... -3. when x = -3 our equation becomes: y=-x-8 y=-(-3)-8 y = 3-8 y = -5

OpenStudy (amistre64):

i dont do essays very well...srry

OpenStudy (amistre64):

with this bounty of knowledge we have just obtained we can use it to make a point on the graph.

OpenStudy (amistre64):

put your crayon on the x = -3 part of the X axis

OpenStudy (anonymous):

Is Mary still there?

OpenStudy (anonymous):

(-3,-5)

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Okay, The drug Lisinopril is used for high blood pressure.

OpenStudy (amistre64):

yes, we put a point at (-3,-5)

OpenStudy (amistre64):

if yall can take this discussion to another post, that would be helpful :)

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

write it in paragraphas

OpenStudy (amistre64):

Now we can use x = 5 and see what we get. y=-x-8 y=-(5)-8 y = -13 (5,-13) gonna have to get longer graph paper for that one ;)

OpenStudy (anonymous):

can we go to the english section and help me i posted there

OpenStudy (anonymous):

graphed

OpenStudy (amistre64):

now we use the other equation, in a different color, and graph its line..

OpenStudy (anonymous):

can we go to the english section and help me i posted there

OpenStudy (amistre64):

9x-8y=-21 , the easiest numbers to use are 0 right?

OpenStudy (anonymous):

right

OpenStudy (amistre64):

ok lets use x = 0 and find a value for y. 9(0)-8y=-21 -8y = -21 y = 21/8 (0,21/8)

OpenStudy (amistre64):

21/8 = 2' 5/8

OpenStudy (anonymous):

amistre64 the essay form is must include : -name of a drug -history of drug -street name -short +long term effect -how is it ingested -what effects does it have on body -anyother intersting facts

OpenStudy (amistre64):

when y = 0 x = what? 9x-8(0)=-21 9x = -21 x = -21/9 (-21/9 , 0) -21/9 = -2' 3/9 = -2' 1/3

OpenStudy (anonymous):

can we go to the english section and help me i posted there

OpenStudy (amistre64):

mary: I dont write essays :) please take it to another post.

OpenStudy (anonymous):

hold on i got lost, im really bad at math

OpenStudy (amistre64):

tell me where we got lost :)

OpenStudy (anonymous):

21/8=2 5/8

OpenStudy (anonymous):

what would my point be for that

OpenStudy (amistre64):

(x=0, y= 2' 5/8) that 'mark helps to define the integer from the fraction

OpenStudy (amistre64):

go up 2 and then mark just a liitle above halfway

OpenStudy (anonymous):

so to graph it (0,2'5/8)

OpenStudy (amistre64):

that would pe one point on the graph of a new line equation.... different colors right?

OpenStudy (anonymous):

right

OpenStudy (amistre64):

the second point is on the x aixs because we made y = 0 right?

OpenStudy (amistre64):

(-2' 1/3 , 0) is our other point, not connect these points with a line

OpenStudy (amistre64):

not means now.... typos persist

OpenStudy (anonymous):

on my paper it asks what is the solution of the equation x=? y=?

OpenStudy (anonymous):

sry solution of the system of equations

OpenStudy (amistre64):

where the lines that we graph cross, where they meet up is a single point. It is asking for the point where the 2 lines meet.

OpenStudy (anonymous):

by my graphing they dont

OpenStudy (amistre64):

we started by finding 2 points for the line to travel thru, but the ends never quit, they go on and on and on.... these lines meet at one point, and that point is our solution. Did you extend your lines until they meet at that point? or did you just connect from dot to dot?

OpenStudy (anonymous):

wait i didnt draw the lines all the way thru

OpenStudy (anonymous):

is it (-5,-2)

OpenStudy (anonymous):

(-5,-3)

OpenStudy (amistre64):

Dunno, you graphed it ;) but lets put those numbers in our equations and see how good they are... y=-x-8 -2 = --5-8 -2 = 5-8 -2 = -3 ..... its a false statment, but we know its close :) 9x-8y=-21 9(-5)-8(-2)=-21 -45 + 16 = -21 -29 = -21 .... wanna try that -3 instead?

OpenStudy (anonymous):

need a ruler, using a folded piece of paper

OpenStudy (amistre64):

the solution by graphing is just a technique, after you use it to "see" what is going on for a solution, you will never really use it again unless you need a visual.

OpenStudy (anonymous):

how often are u on here i need a ft tutor

OpenStudy (amistre64):

as long as the library is open and I dont have class :) ....alot ;)

OpenStudy (anonymous):

ok well i go to strayer university i am always on line, and like i said i really really suck at this kind of math so i always need help got to go for now but will be back on later up until about 3am nc time

OpenStudy (amistre64):

nc ...im in florida :) might be here between 9 and 2

OpenStudy (anonymous):

morning i assume?

OpenStudy (anonymous):

same time zone,

OpenStudy (amistre64):

yep, same zone, college library closes at 2am... I should be around here somewhere. Im the guy with all the fans lol

OpenStudy (anonymous):

yes u r we could exchange numbers if you want that way hearing it sometimes makes better sense

OpenStudy (amistre64):

id rather keep it online ;) you can emailme at tony031172@gmail.com if you need to get in touch with me.

OpenStudy (anonymous):

ok great aatinabo@yahoo